OpenStudy (anonymous):

prime question. f(x)=sqrt(4x) f`(x)=

OpenStudy (harsimran_hs4):

f(x) = sqrt(4x) f'(x) = 1/2 (1/sqrt(4x)) * d/dx (4x) f'(x) = 1/2 (1/sqrt(4x)) * (4)

OpenStudy (anonymous):

so, 4x^1/2?

OpenStudy (anonymous):

I dont understand the way you write.

OpenStudy (harsimran_hs4):

do you know chain rule??

OpenStudy (anonymous):

yes, but im not sure how do i solve prime when sqrtroot is invloved

OpenStudy (anonymous):

I thought the answer is 2/x^-(1/2)?

OpenStudy (harsimran_hs4):

derivative of x^0.5 is 0.5 (x^(0.5-1) = 0.5 (x^(-0.5)) = 0.5 ( 1/x^(0.5))

OpenStudy (anonymous):

mmmm...

OpenStudy (anonymous):

so its 2/x^-(0.5)?

OpenStudy (harsimran_hs4):

what the final answer??

OpenStudy (anonymous):

2/x^-(0.5)?

OpenStudy (harsimran_hs4):

2/ sqrt(4x) 4x not x in your form of answer 2/(4x)^-(0.5)

OpenStudy (anonymous):

Ahhh :( why im getting the wrong answer :/

OpenStudy (anonymous):

I entered the answer and it was wrong

OpenStudy (harsimran_hs4):

well i suggest that you use the equation generator to give the question because otherwise there can be misunderstanding.

OpenStudy (harsimran_hs4):

as i get it it should be \[\frac{ 2 }{ \sqrt{4x} } = \frac{ 1 }{ \sqrt{x} }\]

OpenStudy (anonymous):

Thank you for helping me

OpenStudy (harsimran_hs4):

:)